OpenStudy (anonymous):

Find the slope-intercept form of an equation of the line perpendicular to the graph of x-3y=5 and passing through (0,6). I had y=3x+6 but it's not a choice.

OpenStudy (solomonzelman):

first find the slope of the perpendicular line, do you know how to do this?

OpenStudy (solomonzelman):

I think this example might help. http://openstudy.com/study#/updates/52697f82e4b040e33d2f88e7

OpenStudy (anonymous):

\[Y=\frac{ 1 }{ 3 }X \frac{ 5 }{ 3 }\]

OpenStudy (anonymous):

\[Y=\frac{ 1 }{ 3 }X+\frac{ 5 }{ 3 }\]

OpenStudy (anonymous):

So, it should be y = -3x+6 because you have to do the multiplicative inverse?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

My choices are; A. y = 1/3x - 2 B. y = -3x + 6 C. y = 1/3x + 2 D. y = 3x - 6

OpenStudy (anonymous):

if you take your equation \[x-3y=5\] then all you need to do with it is solve for Y

OpenStudy (solomonzelman):

My bad it is y= 1/3x-3/5 in slope intercept form

OpenStudy (solomonzelman):

So we know that it is going to be y=-3x+b since any perpendicular slope, for example for slope m, is -1/m. now, plug in (0,6) into y=-3x+b to find the missing y-intercept 6=-3(0)+b b=6 so it is going to be y=-3x+6

OpenStudy (anonymous):

Thank you. :D

OpenStudy (solomonzelman):

Anytime